1. cube root of (5/8)
2. sqrt of (25/x^2)
3. sqrt of (x^29)
4. sqrt of (72x^3)

Answer by checkley77(12569)

1. cube root of (5/8)= 5/1.71 or .855 ans.
2. sqrt of (25/x^2)=5/x ans.
3. sqrt of (x^29)x^14(sqrtx) ans
4. sqrt of (72x^3)6x(sqrt2x) ans.